OpenStudy (pinknabastak):

write an equation in standard form of a parabola with zero's of -2 and 5, passing through (1,8)

OpenStudy (zzr0ck3r):

so you know it will sort of look like this (x-5)(x+2) notice if i put in -2, or 5 we get 0 now we take care of the point (1,8) so a(x-5)(x+2) is going to be our form so a(1-5)(1+2) = 8 a(-4)(3) = 8 a(-12)=8 a = -2/3 so our equation is \[\frac{-2}{3}(x-5)(x+2)\]

OpenStudy (pinknabastak):

thanks :)

OpenStudy (zzr0ck3r):

np